OpenStudy (anonymous):

Evaluate 150^2 - 50^2

OpenStudy (anonymous):

20000 :)

OpenStudy (lgbasallote):

you do know that's calculator-able right?

OpenStudy (diyadiya):

\[\Large a^2 -b^2 = (a-b)(a+b)\]\[\Large (150-50)(150+50)\]\[(100 )(200)= ?\]